Day 05-Jul-1999 Shift Report

Arun Tripathi (arun@physics.ucla.edu)
Mon, 5 Jul 1999 13:34:12 -0700 (PDT)


Shift Summary for Day shift, 05/Jul/99

On Shift: Herman White, Arun Tripathi

Hours of beam available at all intensities: 06 hr, 45 min,
Hours of beam available at nominal intensity: ~06 hr, 40 min,
Hours of E832 data (special runs not included): 06 hr, 22 min,
Hours of special runs (laser scans, muon runs...):00 hr

Accounting of lost time:
Unloading tapes from Run 13816 and loading new tapes for run 13817,
initing a new run.

MI RF problems. Beam down for 1 hour and 15 minutes.

========================================================================
Runs ended this shift:
Run 13816 - 24.2 M events - BEAM_832

Tapes brought to FCC this shift:
QKB031, QKC086, QKC087, QKC088, QKK086, QKK087, QKK088,
QKS086, QKS087, QKS088,

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
E832 running.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13816 859 7:34 SEM=1.4E12. Called MCR, and heard someone say
"holy cow" in the background. They will look into
it.
864 7:40 SEM=4.4E12. Tapes fill up, run 13816 stops.
7:46 New CsI calibration constants installed by Ed
(remotely).
13817 7:59 BEAM_832 run 13817 started.
10:00 Marj comes in and notices that alcohol temp. is
-3.7 C. She put two bags of ice around alcohol
refrigerator to see if helps cool it down.
10:55 Alcohol temp. is now -4 C. Ice seems to be
helping.
11:13 He tube #29 opened to renew supply gas pressure,
which was nearly at zero pressure.
12:44 SY energy conservation in effect. No beam.
12:51 MI RF problems. No est.
14:00 Called MCR. Est. 30 min.
14:00 Marj went out and bought some freeze bags (blue
ice). They are being frozen in the KTeV freezer,
to be used for cooling the alcohol refrigerator.
~14:05 Ice around the alcohol frig is gone. Marj went out
and bought some more ice, and put it around the
alcohol bubler.
14:25 Beam comes back.
15:10 Beam is back.
15:30 Hand over to the evening shift.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

In the freezer of the KTeV frig, now there are freeze packs (plastic
boxes with blue stuff inside them). Monitor the alcohol temperature,
especially during the day time, and if the alcohol temp. starts to go up,
put some of these freeze packs around the alcohol frig., in good thermal
contact with it. Do not block the air intake flow on the box. Put any old
freeze pack that has already warmed up back
in the freezer.


[End of Shift Report]